$SPAD/src/input cmds.input

Timothy Daly December 3, 2016

Abstract

Contents

2

— * —

)set break resume
)spool cmds.output
)set message test on
)set message auto off
)clear all
--S 1 of 20
)abbreviation domain TIM TimDaly )quiet
--R
--E 1
--S 2 of 20
)abbreviation domain TIMD TimDalyDomain
--R
--R   TIMD abbreviates domain TimDalyDomain
--E 2
--S 3 of 20
)abbreviation category TIMC TimDalyCategory
--R
--R   TIMC abbreviates category TimDalyCategory
--E 3
--S 4 of 20
)abbreviation package TIMP TimDalyPackage
--R
--R   TIMP abbreviates package TimDalyPackage
--E 4
--S 5 of 20
)abbreviation query LIST
--R
--R   LIST abbreviates domain List
--E 5
--S 6 of 20
)abbreviation query List
--R
--R   LIST abbreviates domain List
--E 6
--S 7 of 20
)abbreviation query TIMD
--R
--R   TIMD abbreviates domain TimDalyDomain
--E 7

--S 8 of 20
)abbreviation remove TIMD
--R
--E 8
--S 9 of 20
)abbreviation query TIMD
--R
--R   TIMD is neither a constructor name nor a constructor abbreviation.
--E 9
--S 10 of 20
)abbreviation query TimDalyPackage
--R
--R   TIMP abbreviates package TimDalyPackage
--E 10
--S 11 of 20
)abbreviation remove TimDalyPackage
--R
--E 11
--S 12 of 20
)abbreviation query TimDalyPackage
--R
--R   TIMP abbreviates package TimDalyPackage
--E 12
--S 13 of 20
)what categories
--R
--R------------------------------- Categories --------------------------------
--R A1AGG    OneDimensionalArrayAggregate ABELGRP  AbelianGroup
--R ABELMON  AbelianMonoid                ABELSG   AbelianSemiGroup
--R ACF      AlgebraicallyClosedField
--R ACFS     AlgebraicallyClosedFunctionSpace
--R AFSPCAT  AffineSpaceCategory          AGG      Aggregate
--R AHYP     ArcHyperbolicFunctionCategory
--R ALAGG    AssociationListAggregate     ALGEBRA  Algebra
--R AMR      AbelianMonoidRing            ARR2CAT  TwoDimensionalArrayCategory
--R ATRIG    ArcTrigonometricFunctionCategory
--R ATTREG   AttributeRegistry            BASTYPE  BasicType
--R BGAGG    BagAggregate                 BLMETCT  BlowUpMethodCategory
--R BMODULE  BiModule                     BRAGG    BinaryRecursiveAggregate
--R BTAGG    BitAggregate                 BTCAT    BinaryTreeCategory
--R CABMON   CancellationAbelianMonoid    CACHSET  CachableSet
--R CFCAT    CombinatorialFunctionCategory
--R CHARNZ   CharacteristicNonZero        CHARZ    CharacteristicZero
--R CLAGG    Collection                   COMBOPC  CombinatorialOpsCategory
--R COMPCAT  ComplexCategory              COMRING  CommutativeRing

--R DIAGG    Dictionary                   DIFEXT   DifferentialExtension
--R DIFRING  DifferentialRing             DIOPS    DictionaryOperations
--R DIRPCAT  DirectProductCategory        DIVCAT   DivisorCategory
--R DIVRING  DivisionRing                 DLAGG    DoublyLinkedAggregate
--R DPOLCAT  DifferentialPolynomialCategory
--R DQAGG    DequeueAggregate             DSTRCAT  DesingTreeCategory
--R DVARCAT  DifferentialVariableCategory ELAGG    ExtensibleLinearAggregate
--R ELEMFUN  ElementaryFunctionCategory   ELTAB    Eltable
--R ELTAGG   EltableAggregate             ENTIRER  EntireRing
--R ES       ExpressionSpace              EUCDOM   EuclideanDomain
--R EVALAB   Evalable                     FAMONC   FreeAbelianMonoidCategory
--R FAMR     FiniteAbelianMonoidRing
--R FAXF     FiniteAlgebraicExtensionField
--R FDIVCAT  FiniteDivisorCategory        FEVALAB  FullyEvalableOver
--R FFCAT    FunctionFieldCategory        FFIELDC  FiniteFieldCategory
--R FIELD    Field                        FILECAT  FileCategory
--R FINAALG  FiniteRankNonAssociativeAlgebra
--R FINITE   Finite                       FINRALG  FiniteRankAlgebra
--R FLAGG    FiniteLinearAggregate        FLALG    FreeLieAlgebra
--R FLINEXP  FullyLinearlyExplicitRingOver
--R FMC      FortranMatrixCategory        FMCAT    FreeModuleCat
--R FMFUN    FortranMatrixFunctionCategory
--R FMTC     FortranMachineTypeCategory   FNCAT    FileNameCategory
--R FORTCAT  FortranProgramCategory       FORTFN   FortranFunctionCategory
--R FPATMAB  FullyPatternMatchable        FPC      FieldOfPrimeCharacteristic
--R FPS      FloatingPointSystem          FRAMALG  FramedAlgebra
--R FRETRCT  FullyRetractableTo           FRNAALG  FramedNonAssociativeAlgebra
--R FS       FunctionSpace                FSAGG    FiniteSetAggregate
--R FVC      FortranVectorCategory
--R FVFUN    FortranVectorFunctionCategory
--R GCDDOM   GcdDomain                    GRALG    GradedAlgebra
--R GRMOD    GradedModule                 GROUP    Group
--R HOAGG    HomogeneousAggregate         HYPCAT   HyperbolicFunctionCategory
--R IDPC     IndexedDirectProductCategory IEVALAB  InnerEvalable
--R INFCLCT  InfinitlyClosePointCategory  INS      IntegerNumberSystem
--R INTCAT   IntervalCategory             INTDOM   IntegralDomain
--R IXAGG    IndexedAggregate             KDAGG    KeyedDictionary
--R KOERCE   CoercibleTo                  KONVERT  ConvertibleTo
--R LALG     LeftAlgebra                  LFCAT    LiouvillianFunctionCategory
--R LIECAT   LieAlgebra                   LINEXP   LinearlyExplicitRingOver
--R LMODULE  LeftModule                   LNAGG    LinearAggregate
--R LOCPOWC  LocalPowerSeriesCategory
--R LODOCAT  LinearOrdinaryDifferentialOperatorCategory
--R LOGIC    Logic                        LSAGG    ListAggregate
--R LZSTAGG  LazyStreamAggregate
--R MAGCDOC  ModularAlgebraicGcdOperations
--R MATCAT   MatrixCategory               MDAGG    MultiDictionary
--R MLO      MonogenicLinearOperator      MODULE   Module
--R MONAD    Monad                        MONADWU  MonadWithUnit
--R MONOGEN  MonogenicAlgebra             MONOID   Monoid

--R MSETAGG  MultisetAggregate
--R MTSCAT   MultivariateTaylorSeriesCategory
--R NAALG    NonAssociativeAlgebra        NARNG    NonAssociativeRng
--R NASRING  NonAssociativeRing
--R NTSCAT   NormalizedTriangularSetCategory
--R NUMINT   NumericalIntegrationCategory OAGROUP  OrderedAbelianGroup
--R OAMON    OrderedAbelianMonoid         OAMONS   OrderedAbelianMonoidSup
--R OASGP    OrderedAbelianSemiGroup      OC       OctonionCategory
--R OCAMON   OrderedCancellationAbelianMonoid
--R ODECAT   OrdinaryDifferentialEquationsSolverCategory
--R OINTDOM  OrderedIntegralDomain        OM       OpenMath
--R OMSAGG   OrderedMultisetAggregate
--R OPTCAT   NumericalOptimizationCategory
--R ORDFIN   OrderedFinite                ORDMON   OrderedMonoid
--R ORDRING  OrderedRing                  ORDSET   OrderedSet
--R OREPCAT  UnivariateSkewPolynomialCategory
--R PACEXTC  PseudoAlgebraicClosureOfAlgExtOfRationalNumberCategory
--R PACFFC   PseudoAlgebraicClosureOfFiniteFieldCategory
--R PACPERC  PseudoAlgebraicClosureOfPerfectFieldCategory
--R PACRATC  PseudoAlgebraicClosureOfRationalNumberCategory
--R PADICCT  PAdicIntegerCategory         PATAB    Patternable
--R PATMAB   PatternMatchable
--R PDECAT   PartialDifferentialEquationsSolverCategory
--R PDRING   PartialDifferentialRing      PERMCAT  PermutationCategory
--R PFECAT   PolynomialFactorizationExplicit
--R PID      PrincipalIdealDomain         PLACESC  PlacesCategory
--R POLYCAT  PolynomialCategory           PPCURVE  PlottablePlaneCurveCategory
--R PRIMCAT  PrimitiveFunctionCategory    PRQAGG   PriorityQueueAggregate
--R PRSPCAT  ProjectiveSpaceCategory      PSCAT    PowerSeriesCategory
--R PSCURVE  PlottableSpaceCurveCategory  PSETCAT  PolynomialSetCategory
--R PTCAT    PointCategory
--R PTRANFN  PartialTranscendentalFunctions
--R QFCAT    QuotientFieldCategory        QUAGG    QueueAggregate
--R QUATCAT  QuaternionCategory           RADCAT   RadicalCategory
--R RCAGG    RecursiveAggregate           RCFIELD  RealClosedField
--R REAL     RealConstant                 RETRACT  RetractableTo
--R RING     Ring                         RMATCAT  RectangularMatrixCategory
--R RMODULE  RightModule                  RNG      Rng
--R RNS      RealNumberSystem             RPOLCAT  RecursivePolynomialCategory
--R RRCC     RealRootCharacterizationCategory
--R RSETCAT  RegularTriangularSetCategory SEGCAT   SegmentCategory
--R SEGXCAT  SegmentExpansionCategory     SETAGG   SetAggregate
--R SETCAT   SetCategory                  SETCATD  SetCategoryWithDegree
--R SEXCAT   SExpressionCategory
--R SFRTCAT  SquareFreeRegularTriangularSetCategory
--R SGROUP   SemiGroup                    SKAGG    StackAggregate
--R SMATCAT  SquareMatrixCategory
--R SNTSCAT  SquareFreeNormalizedTriangularSetCategory
--R SPACEC   ThreeSpaceCategory           SPFCAT   SpecialFunctionCategory
--R SRAGG    StringAggregate              STAGG    StreamAggregate

--R STEP     StepThrough                  STRICAT  StringCategory
--R TBAGG    TableAggregate
--R TRANFUN  TranscendentalFunctionCategory
--R TRIGCAT  TrigonometricFunctionCategory
--R TSETCAT  TriangularSetCategory        TYPE     Type
--R UFD      UniqueFactorizationDomain
--R ULSCAT   UnivariateLaurentSeriesCategory
--R ULSCCAT  UnivariateLaurentSeriesConstructorCategory
--R UPOLYC   UnivariatePolynomialCategory
--R UPSCAT   UnivariatePowerSeriesCategory
--R UPXSCAT  UnivariatePuiseuxSeriesCategory
--R UPXSCCA  UnivariatePuiseuxSeriesConstructorCategory
--R URAGG    UnaryRecursiveAggregate
--R UTSCAT   UnivariateTaylorSeriesCategory
--R VECTCAT  VectorCategory               VSPACE   VectorSpace
--R XALG     XAlgebra                     XF       ExtensionField
--R XFALG    XFreeAlgebra                 XPOLYC   XPolynomialsCat
--E 13
--S 14 of 20
)what domains
--R
--R--------------------------------- Domains ---------------------------------
--R A1AGG-   OneDimensionalArrayAggregate&
--R ABELGRP- AbelianGroup&                ABELMON- AbelianMonoid&
--R ABELSG-  AbelianSemiGroup&            ACF-     AlgebraicallyClosedField&
--R ACFS-    AlgebraicallyClosedFunctionSpace&
--R ACPLOT   PlaneAlgebraicCurvePlot      AFFPL    AffinePlane
--R AFFPLPS  AffinePlaneOverPseudoAlgebraicClosureOfFiniteField
--R AFFSP    AffineSpace                  AGG-     Aggregate&
--R ALGEBRA- Algebra&                     ALGFF    AlgebraicFunctionField
--R ALGSC    AlgebraGivenByStructuralConstants
--R ALIST    AssociationList              AMR-     AbelianMonoidRing&
--R AN       AlgebraicNumber              ANON     AnonymousFunction
--R ANTISYM  AntiSymm                     ANY      Any
--R ARR2CAT- TwoDimensionalArrayCategory& ARRAY1   OneDimensionalArray
--R ARRAY2   TwoDimensionalArray          ASP1     Asp1
--R ASP10    Asp10                        ASP12    Asp12
--R ASP19    Asp19                        ASP20    Asp20
--R ASP24    Asp24                        ASP27    Asp27
--R ASP28    Asp28                        ASP29    Asp29
--R ASP30    Asp30                        ASP31    Asp31
--R ASP33    Asp33                        ASP34    Asp34
--R ASP35    Asp35                        ASP4     Asp4
--R ASP41    Asp41                        ASP42    Asp42
--R ASP49    Asp49                        ASP50    Asp50
--R ASP55    Asp55                        ASP6     Asp6
--R ASP7     Asp7                         ASP73    Asp73
--R ASP74    Asp74                        ASP77    Asp77
--R ASP78    Asp78                        ASP8     Asp8

--R ASP80    Asp80                        ASP9     Asp9
--R ASTACK   ArrayStack
--R ATRIG-   ArcTrigonometricFunctionCategory&
--R ATTRBUT  AttributeButtons             AUTOMOR  Automorphism
--R BASTYPE- BasicType&                   BBTREE   BalancedBinaryTree
--R BFUNCT   BasicFunctions               BGAGG-   BagAggregate&
--R BINARY   BinaryExpansion              BINFILE  BinaryFile
--R BITS     Bits                         BLHN     BlowUpWithHamburgerNoether
--R BLQT     BlowUpWithQuadTrans          BOOLEAN  Boolean
--R BOP      BasicOperator                BPADIC   BalancedPAdicInteger
--R BPADICRT BalancedPAdicRational        BRAGG-   BinaryRecursiveAggregate&
--R BSD      BasicStochasticDifferential  BSTREE   BinarySearchTree
--R BTAGG-   BitAggregate&                BTCAT-   BinaryTreeCategory&
--R BTOURN   BinaryTournament             BTREE    BinaryTree
--R CARD     CardinalNumber               CARTEN   CartesianTensor
--R CCLASS   CharacterClass               CDFMAT   ComplexDoubleFloatMatrix
--R CDFVEC   ComplexDoubleFloatVector     CHAR     Character
--R CLAGG-   Collection&                  CLIF     CliffordAlgebra
--R COLOR    Color                        COMM     Commutator
--R COMPCAT- ComplexCategory&             COMPLEX  Complex
--R COMPPROP SubSpaceComponentProperty    CONTFRAC ContinuedFraction
--R D01AJFA  d01ajfAnnaType               D01AKFA  d01akfAnnaType
--R D01ALFA  d01alfAnnaType               D01AMFA  d01amfAnnaType
--R D01ANFA  d01anfAnnaType               D01APFA  d01apfAnnaType
--R D01AQFA  d01aqfAnnaType               D01ASFA  d01asfAnnaType
--R D01FCFA  d01fcfAnnaType               D01GBFA  d01gbfAnnaType
--R D01TRNS  d01TransformFunctionType     D02BBFA  d02bbfAnnaType
--R D02BHFA  d02bhfAnnaType               D02CJFA  d02cjfAnnaType
--R D02EJFA  d02ejfAnnaType               D03EEFA  d03eefAnnaType
--R D03FAFA  d03fafAnnaType               DBASE    Database
--R DECIMAL  DecimalExpansion             DEQUEUE  Dequeue
--R DERHAM   DeRhamComplex                DFLOAT   DoubleFloat
--R DFMAT    DoubleFloatMatrix            DFVEC    DoubleFloatVector
--R DHMATRIX DenavitHartenbergMatrix      DIAGG-   Dictionary&
--R DIFEXT-  DifferentialExtension&       DIFRING- DifferentialRing&
--R DIOPS-   DictionaryOperations&        DIRPCAT- DirectProductCategory&
--R DIRPROD  DirectProduct                DIRRING  DirichletRing
--R DIV      Divisor                      DIVRING- DivisionRing&
--R DLIST    DataList
--R DMP      DistributedMultivariatePolynomial
--R DPMM     DirectProductMatrixModule    DPMO     DirectProductModule
--R DPOLCAT- DifferentialPolynomialCategory&
--R DROPT    DrawOption
--R DSMP     DifferentialSparseMultivariatePolynomial
--R DSTREE   DesingTree
--R DVARCAT- DifferentialVariableCategory&
--R E04DGFA  e04dgfAnnaType               E04FDFA  e04fdfAnnaType
--R E04GCFA  e04gcfAnnaType               E04JAFA  e04jafAnnaType
--R E04MBFA  e04mbfAnnaType               E04NAFA  e04nafAnnaType
--R E04UCFA  e04ucfAnnaType               EAB      ExtAlgBasis

--R EFULS    ElementaryFunctionsUnivariateLaurentSeries
--R EFUPXS   ElementaryFunctionsUnivariatePuiseuxSeries
--R ELAGG-   ExtensibleLinearAggregate&   ELEMFUN- ElementaryFunctionCategory&
--R ELTAGG-  EltableAggregate&            EMR      EuclideanModularRing
--R EQ       Equation                     EQTBL    EqTable
--R ES-      ExpressionSpace&             EUCDOM-  EuclideanDomain&
--R EVALAB-  Evalable&                    EXIT     Exit
--R EXPEXPAN ExponentialExpansion         EXPR     Expression
--R EXPUPXS  ExponentialOfUnivariatePuiseuxSeries
--R FAGROUP  FreeAbelianGroup             FAMONOID FreeAbelianMonoid
--R FAMR-    FiniteAbelianMonoidRing&     FARRAY   FlexibleArray
--R FAXF-    FiniteAlgebraicExtensionField&
--R FC       FortranCode                  FCOMP    FourierComponent
--R FDIV     FiniteDivisor                FDIVCAT- FiniteDivisorCategory&
--R FEVALAB- FullyEvalableOver&           FEXPR    FortranExpression
--R FF       FiniteField                  FFCAT-   FunctionFieldCategory&
--R FFCG     FiniteFieldCyclicGroup
--R FFCGP    FiniteFieldCyclicGroupExtensionByPolynomial
--R FFCGX    FiniteFieldCyclicGroupExtension
--R FFIELDC- FiniteFieldCategory&         FFNB     FiniteFieldNormalBasis
--R FFNBP    FiniteFieldNormalBasisExtensionByPolynomial
--R FFNBX    FiniteFieldNormalBasisExtension
--R FFP      FiniteFieldExtensionByPolynomial
--R FFX      FiniteFieldExtension         FGROUP   FreeGroup
--R FIELD-   Field&                       FILE     File
--R FINAALG- FiniteRankNonAssociativeAlgebra&
--R FINRALG- FiniteRankAlgebra&           FLAGG-   FiniteLinearAggregate&
--R FLINEXP- FullyLinearlyExplicitRingOver&
--R FLOAT    Float                        FM       FreeModule
--R FM1      FreeModule1                  FMONOID  FreeMonoid
--R FNAME    FileName                     FNLA     FreeNilpotentLie
--R FORMULA  ScriptFormulaFormat          FORTRAN  FortranProgram
--R FPARFRAC FullPartialFractionExpansion FPC-     FieldOfPrimeCharacteristic&
--R FPS-     FloatingPointSystem&         FR       Factored
--R FRAC     Fraction                     FRAMALG- FramedAlgebra&
--R FRETRCT- FullyRetractableTo&          FRIDEAL  FractionalIdeal
--R FRMOD    FramedModule                 FRNAALG- FramedNonAssociativeAlgebra&
--R FS-      FunctionSpace&               FSAGG-   FiniteSetAggregate&
--R FSERIES  FourierSeries                FST      FortranScalarType
--R FT       FortranType                  FTEM     FortranTemplate
--R FUNCTION FunctionCalled               GCDDOM-  GcdDomain&
--R GCNAALG  GenericNonAssociativeAlgebra
--R GDMP     GeneralDistributedMultivariatePolynomial
--R GMODPOL  GeneralModulePolynomial      GOPT     GuessOption
--R GOPT0    GuessOptionFunctions0        GPOLSET  GeneralPolynomialSet
--R GRALG-   GradedAlgebra&               GRIMAGE  GraphImage
--R GRMOD-   GradedModule&                GROUP-   Group&
--R GSERIES  GeneralUnivariatePowerSeries GSTBL    GeneralSparseTable
--R GTSET    GeneralTriangularSet         HACKPI   Pi
--R HASHTBL  HashTable

--R HDMP     HomogeneousDistributedMultivariatePolynomial
--R HDP      HomogeneousDirectProduct     HEAP     Heap
--R HELLFDIV HyperellipticFiniteDivisor   HEXADEC  HexadecimalExpansion
--R HOAGG-   HomogeneousAggregate&        HTMLFORM HTMLFormat
--R HYPCAT-  HyperbolicFunctionCategory&  IAN      InnerAlgebraicNumber
--R IARRAY1  IndexedOneDimensionalArray   IARRAY2  IndexedTwoDimensionalArray
--R IBITS    IndexedBits                  ICARD    IndexCard
--R ICP      InfClsPt                     IDEAL    PolynomialIdeals
--R IDPAG    IndexedDirectProductAbelianGroup
--R IDPAM    IndexedDirectProductAbelianMonoid
--R IDPO     IndexedDirectProductObject
--R IDPOAM   IndexedDirectProductOrderedAbelianMonoid
--R IDPOAMS  IndexedDirectProductOrderedAbelianMonoidSup
--R IEVALAB- InnerEvalable&               IFAMON   InnerFreeAbelianMonoid
--R IFARRAY  IndexedFlexibleArray         IFF      InnerFiniteField
--R IIARRAY2 InnerIndexedTwoDimensionalArray
--R ILIST    IndexedList                  IMATRIX  IndexedMatrix
--R INDE     IndexedExponents
--R INFCLSPS InfinitlyClosePointOverPseudoAlgebraicClosureOfFiniteField
--R INFCLSPT InfinitlyClosePoint          INFORM   InputForm
--R INS-     IntegerNumberSystem&         INT      Integer
--R INTABL   InnerTable                   INTDOM-  IntegralDomain&
--R INTFTBL  IntegrationFunctionsTable    INTRVL   Interval
--R IPADIC   InnerPAdicInteger            IPF      InnerPrimeField
--R IR       IntegrationResult            ISTRING  IndexedString
--R ISUPS    InnerSparseUnivariatePowerSeries
--R ITAYLOR  InnerTaylorSeries            ITUPLE   InfiniteTuple
--R IVECTOR  IndexedVector                IXAGG-   IndexedAggregate&
--R JORDAN   AssociatedJordanAlgebra      KAFILE   KeyedAccessFile
--R KDAGG-   KeyedDictionary&             KERNEL   Kernel
--R LA       LocalAlgebra                 LALG-    LeftAlgebra&
--R LAUPOL   LaurentPolynomial            LEXP     LieExponentials
--R LIB      Library                      LIE      AssociatedLieAlgebra
--R LIECAT-  LieAlgebra&                  LIST     List
--R LMDICT   ListMultiDictionary          LMOPS    ListMonoidOps
--R LNAGG-   LinearAggregate&             LO       Localize
--R LODO     LinearOrdinaryDifferentialOperator
--R LODO1    LinearOrdinaryDifferentialOperator1
--R LODO2    LinearOrdinaryDifferentialOperator2
--R LODOCAT- LinearOrdinaryDifferentialOperatorCategory&
--R LOGIC-   Logic&                       LPOLY    LiePolynomial
--R LSAGG-   ListAggregate&               LSQM     LieSquareMatrix
--R LWORD    LyndonWord                   LZSTAGG- LazyStreamAggregate&
--R M3D      ThreeDimensionalMatrix       MAGMA    Magma
--R MATCAT-  MatrixCategory&              MATRIX   Matrix
--R MCMPLX   MachineComplex               MFLOAT   MachineFloat
--R MINT     MachineInteger               MKCHSET  MakeCachableSet
--R MMLFORM  MathMLFormat                 MODFIELD ModularField
--R MODMON   ModMonic                     MODMONOM ModuleMonomial
--R MODOP    ModuleOperator               MODRING  ModularRing

--R MODULE-  Module&                      MOEBIUS  MoebiusTransform
--R MONAD-   Monad&                       MONADWU- MonadWithUnit&
--R MONOGEN- MonogenicAlgebra&            MONOID-  Monoid&
--R MPOLY    MultivariatePolynomial       MRING    MonoidRing
--R MSET     Multiset                     MYEXPR   MyExpression
--R MYUP     MyUnivariatePolynomial       NAALG-   NonAssociativeAlgebra&
--R NARNG-   NonAssociativeRng&           NASRING- NonAssociativeRing&
--R       Enumeration                           Mapping
--R       Record                                Union
--R NIPROB   NumericalIntegrationProblem  NNI      NonNegativeInteger
--R NONE     None                         NOTTING  NottinghamGroup
--R NSDPS    NeitherSparseOrDensePowerSeries
--R NSMP     NewSparseMultivariatePolynomial
--R NSUP     NewSparseUnivariatePolynomial
--R OC-      OctonionCategory&            OCT      Octonion
--R ODEIFTBL ODEIntensityFunctionsTable   ODEPROB  NumericalODEProblem
--R ODP      OrderedDirectProduct
--R ODPOL    OrderlyDifferentialPolynomial
--R ODR      OrdinaryDifferentialRing     ODVAR    OrderlyDifferentialVariable
--R OFMONOID OrderedFreeMonoid            OMCONN   OpenMathConnection
--R OMDEV    OpenMathDevice               OMENC    OpenMathEncoding
--R OMERR    OpenMathError                OMERRK   OpenMathErrorKind
--R OMLO     OppositeMonogenicLinearOperator
--R ONECOMP  OnePointCompletion           OP       Operator
--R OPTPROB  NumericalOptimizationProblem ORDCOMP  OrderedCompletion
--R ORDRING- OrderedRing&                 ORDSET-  OrderedSet&
--R OREPCAT- UnivariateSkewPolynomialCategory&
--R ORESUP   SparseUnivariateSkewPolynomial
--R OREUP    UnivariateSkewPolynomial     OSI      OrdSetInts
--R OUTFORM  OutputForm                   OVAR     OrderedVariableList
--R OWP      OrdinaryWeightedPolynomials
--R PACEXT   PseudoAlgebraicClosureOfAlgExtOfRationalNumber
--R PACOFF   PseudoAlgebraicClosureOfFiniteField
--R PACRAT   PseudoAlgebraicClosureOfRationalNumber
--R PADIC    PAdicInteger                 PADICRAT PAdicRational
--R PADICRC  PAdicRationalConstructor     PALETTE  Palette
--R PARPCURV ParametricPlaneCurve         PARSCURV ParametricSpaceCurve
--R PARSURF  ParametricSurface            PATLRES  PatternMatchListResult
--R PATRES   PatternMatchResult           PATTERN  Pattern
--R PBWLB    PoincareBirkhoffWittLyndonBasis
--R PDEPROB  NumericalPDEProblem          PDRING-  PartialDifferentialRing&
--R PENDTREE PendantTree                  PERM     Permutation
--R PERMGRP  PermutationGroup             PF       PrimeField
--R PFECAT-  PolynomialFactorizationExplicit&
--R PFR      PartialFraction              PI       PositiveInteger
--R PLACES   Places
--R PLACESPS PlacesOverPseudoAlgebraicClosureOfFiniteField
--R PLCS     Plcs                         PLOT     Plot
--R PLOT3D   Plot3D                       POINT    Point
--R POLY     Polynomial                   POLYCAT- PolynomialCategory&

--R PR       PolynomialRing               PRIMARR  PrimitiveArray
--R PRODUCT  Product                      PROJPL   ProjectivePlane
--R PROJPLPS ProjectivePlaneOverPseudoAlgebraicClosureOfFiniteField
--R PROJSP   ProjectiveSpace              PRTITION Partition
--R PSCAT-   PowerSeriesCategory&         PSETCAT- PolynomialSetCategory&
--R QALGSET  QuasiAlgebraicSet            QEQUAT   QueryEquation
--R QFCAT-   QuotientFieldCategory&       QFORM    QuadraticForm
--R QUAT     Quaternion                   QUATCAT- QuaternionCategory&
--R QUEUE    Queue                        RADCAT-  RadicalCategory&
--R RADFF    RadicalFunctionField         RADIX    RadixExpansion
--R RCAGG-   RecursiveAggregate&          RCFIELD- RealClosedField&
--R RECLOS   RealClosure                  REF      Reference
--R REGSET   RegularTriangularSet         RESRING  ResidueRing
--R RESULT   Result                       RETRACT- RetractableTo&
--R RGCHAIN  RegularChain                 RING-    Ring&
--R RMATCAT- RectangularMatrixCategory&   RMATRIX  RectangularMatrix
--R RNS-     RealNumberSystem&
--R ROIRC    RightOpenIntervalRootCharacterization
--R ROMAN    RomanNumeral                 ROUTINE  RoutinesTable
--R RPOLCAT- RecursivePolynomialCategory&
--R RRCC-    RealRootCharacterizationCategory&
--R RSETCAT- RegularTriangularSetCategory&
--R RULE     RewriteRule                  RULECOLD RuleCalled
--R RULESET  Ruleset                      SAE      SimpleAlgebraicExtension
--R SAOS     SingletonAsOrderedSet        SD       StochasticDifferential
--R SDPOL    SequentialDifferentialPolynomial
--R SDVAR    SequentialDifferentialVariable
--R SEG      Segment                      SEGBIND  SegmentBinding
--R SET      Set                          SETAGG-  SetAggregate&
--R SETCAT-  SetCategory&                 SETMN    SetOfMIntegersInOneToN
--R SEX      SExpression                  SEXOF    SExpressionOf
--R SFORT    SimpleFortranProgram         SGROUP-  SemiGroup&
--R SHDP     SplitHomogeneousDirectProduct
--R SINT     SingleInteger                SMATCAT- SquareMatrixCategory&
--R SMP      SparseMultivariatePolynomial
--R SMTS     SparseMultivariateTaylorSeries
--R SPACE3   ThreeSpace                   SPLNODE  SplittingNode
--R SPLTREE  SplittingTree                SQMATRIX SquareMatrix
--R SRAGG-   StringAggregate&
--R SREGSET  SquareFreeRegularTriangularSet
--R STACK    Stack                        STAGG-   StreamAggregate&
--R STBL     SparseTable                  STREAM   Stream
--R STRING   String                       STRTBL   StringTable
--R SUBSPACE SubSpace                     SUCH     SuchThat
--R SULS     SparseUnivariateLaurentSeries
--R SUP      SparseUnivariatePolynomial
--R SUPEXPR  SparseUnivariatePolynomialExpressions
--R SUPXS    SparseUnivariatePuiseuxSeries
--R SUTS     SparseUnivariateTaylorSeries SWITCH   Switch
--R SYMBOL   Symbol                       SYMPOLY  SymmetricPolynomial

--R SYMS     TheSymbolTable               SYMTAB   SymbolTable
--R TABLE    Table                        TABLEAU  Tableau
--R TBAGG-   TableAggregate&              TEX      TexFormat
--R TEXTFILE TextFile
--R TRANFUN- TranscendentalFunctionCategory&
--R TREE     Tree
--R TRIGCAT- TrigonometricFunctionCategory&
--R TS       TaylorSeries                 TSETCAT- TriangularSetCategory&
--R TUBE     TubePlot                     TUPLE    Tuple
--R U16MAT   U16Matrix                    U16VEC   U16Vector
--R U32MAT   U32Matrix                    U32VEC   U32Vector
--R U8MAT    U8Matrix                     U8VEC    U8Vector
--R UFD-     UniqueFactorizationDomain&   UFPS     UnivariateFormalPowerSeries
--R ULS      UnivariateLaurentSeries
--R ULSCCAT- UnivariateLaurentSeriesConstructorCategory&
--R ULSCONS  UnivariateLaurentSeriesConstructor
--R UNISEG   UniversalSegment             UP       UnivariatePolynomial
--R UPOLYC-  UnivariatePolynomialCategory&
--R UPSCAT-  UnivariatePowerSeriesCategory&
--R UPXS     UnivariatePuiseuxSeries
--R UPXSCCA- UnivariatePuiseuxSeriesConstructorCategory&
--R UPXSCONS UnivariatePuiseuxSeriesConstructor
--R UPXSSING UnivariatePuiseuxSeriesWithExponentialSingularity
--R URAGG-   UnaryRecursiveAggregate&     UTS      UnivariateTaylorSeries
--R UTSCAT-  UnivariateTaylorSeriesCategory&
--R UTSZ     UnivariateTaylorSeriesCZero  VARIABLE Variable
--R VECTCAT- VectorCategory&              VECTOR   Vector
--R VIEW2D   TwoDimensionalViewport       VIEW3D   ThreeDimensionalViewport
--R VOID     Void                         VSPACE-  VectorSpace&
--R WP       WeightedPolynomials          WUTSET   WuWenTsunTriangularSet
--R XDPOLY   XDistributedPolynomial       XF-      ExtensionField&
--R XPBWPOLY XPBWPolynomial               XPOLY    XPolynomial
--R XPR      XPolynomialRing              XRPOLY   XRecursivePolynomial
--R ZMOD     IntegerMod
--E 14
--S 15 of 20
)what packages
--R
--R-------------------------------- Packages ---------------------------------
--R AF       AlgebraicFunction
--R AFALGGRO AffineAlgebraicSetComputeWithGroebnerBasis
--R AFALGRES AffineAlgebraicSetComputeWithResultant
--R ALGFACT  AlgFactor                    ALGMANIP AlgebraicManipulations
--R ALGMFACT AlgebraicMultFact            ALGPKG   AlgebraPackage
--R ANY1     AnyFunctions1                API      ApplicationProgramInterface
--R APPLYORE ApplyUnivariateSkewPolynomial
--R APPRULE  ApplyRules
--R ARRAY12  OneDimensionalArrayFunctions2
--R ASSOCEQ  AssociatedEquations          AXSERV   AxiomServer

--R BALFACT  BalancedFactorisation        BEZIER   Bezier
--R BEZOUT   BezoutMatrix                 BLAS1    BlasLevelOne
--R BLUPPACK BlowUpPackage                BOP1     BasicOperatorFunctions1
--R BOUNDZRO BoundIntegerRoots            BRILL    BrillhartTests
--R CARTEN2  CartesianTensorFunctions2    CDEN     CommonDenominator
--R CHARPOL  CharacteristicPolynomialPackage
--R CHVAR    ChangeOfVariable
--R CINTSLPE ComplexIntegerSolveLinearPolynomialEquation
--R CLIP     TwoDimensionalPlotClipping   CMPLXRT  ComplexRootPackage
--R COMBF    CombinatorialFunction        COMBINAT IntegerCombinatoricFunctions
--R COMMONOP CommonOperators
--R COMMUPC  CommuteUnivariatePolynomialCategory
--R COMPFACT ComplexFactorization         COMPLEX2 ComplexFunctions2
--R COMPLPAT ComplexPattern               COORDSYS CoordinateSystems
--R CPIMA    CharacteristicPolynomialInMonogenicalAlgebra
--R CPMATCH  ComplexPatternMatch          CRAPACK  CRApackage
--R CRFP     ComplexRootFindingPackage    CSTTOOLS CyclicStreamTools
--R CTRIGMNP ComplexTrigonometricManipulations
--R CVMP     CoerceVectorMatrixPackage    CYCLES   CycleIndicators
--R CYCLOTOM CyclotomicPolynomialPackage  D01AGNT  d01AgentsPackage
--R D01WGTS  d01WeightsPackage            D02AGNT  d02AgentsPackage
--R D03AGNT  d03AgentsPackage             DBLRESP  DoubleResultantPackage
--R DDFACT   DistinctDegreeFactorize
--R DEFINTEF ElementaryFunctionDefiniteIntegration
--R DEFINTRF RationalFunctionDefiniteIntegration
--R DEGRED   DegreeReductionPackage       DFINTTLS DefiniteIntegrationTools
--R DFSFUN   DoubleFloatSpecialFunctions  DIOSP    DiophantineSolutionPackage
--R DIRPROD2 DirectProductFunctions2      DISPLAY  DisplayPackage
--R DLP      DiscreteLogarithmPackage     DRAW     TopLevelDrawFunctions
--R DRAWCFUN TopLevelDrawFunctionsForCompiledFunctions
--R DRAWCURV TopLevelDrawFunctionsForAlgebraicCurves
--R DRAWCX   DrawComplex                  DRAWHACK DrawNumericHack
--R DRAWPT   TopLevelDrawFunctionsForPoints
--R DROPT0   DrawOptionFunctions0         DROPT1   DrawOptionFunctions1
--R DTP      DesingTreePackage            E04AGNT  e04AgentsPackage
--R EF       ElementaryFunction
--R EFSTRUC  ElementaryFunctionStructurePackage
--R ELFUTS   EllipticFunctionsUnivariateTaylorSeries
--R EP       EigenPackage                 EQ2      EquationFunctions2
--R ERROR    ErrorFunctions               ES1      ExpressionSpaceFunctions1
--R ES2      ExpressionSpaceFunctions2
--R ESCONT   ExpertSystemContinuityPackage
--R ESCONT1  ExpertSystemContinuityPackage1
--R ESTOOLS  ExpertSystemToolsPackage     ESTOOLS1 ExpertSystemToolsPackage1
--R ESTOOLS2 ExpertSystemToolsPackage2    EVALCYC  EvaluateCycleIndicators
--R EXP3D    Export3D                     EXPR2    ExpressionFunctions2
--R EXPR2UPS ExpressionToUnivariatePowerSeries
--R EXPRODE  ExpressionSpaceODESolver     EXPRSOL  ExpressionSolve
--R EXPRTUBE ExpressionTubePlot
--R FACTEXT  FactorisationOverPseudoAlgebraicClosureOfAlgExtOfRationalNumber

--R FACTFUNC FactoredFunctions
--R FACTRN   FactorisationOverPseudoAlgebraicClosureOfRationalNumber
--R FACUTIL  FactoringUtilities
--R FAMR2    FiniteAbelianMonoidRingFunctions2
--R FCPAK1   FortranCodePackage1          FDIV2    FiniteDivisorFunctions2
--R FFCAT2   FunctionFieldCategoryFunctions2
--R FFF      FiniteFieldFunctions         FFFACTOR FiniteFieldFactorization
--R FFFACTSE FiniteFieldFactorizationWithSizeParseBySideEffect
--R FFFG     FractionFreeFastGaussian
--R FFFGF    FractionFreeFastGaussianFractions
--R FFHOM    FiniteFieldHomomorphisms     FFINTBAS FunctionFieldIntegralBasis
--R FFPOLY   FiniteFieldPolynomialPackage
--R FFPOLY2  FiniteFieldPolynomialPackage2
--R FFSLPE   FiniteFieldSolveLinearPolynomialEquation
--R FFSQFR   FiniteFieldSquareFreeDecomposition
--R FGLMICPK FGLMIfCanPackage
--R FLAGG2   FiniteLinearAggregateFunctions2
--R FLASORT  FiniteLinearAggregateSort    FLOATCP  FloatingComplexPackage
--R FLOATRP  FloatingRealPackage          FOP      FortranOutputStackPackage
--R FORDER   FindOrderFinite              FORMULA1 ScriptFormulaFormat1
--R FORT     FortranPackage               FR2      FactoredFunctions2
--R FRAC2    FractionFunctions2           FRIDEAL2 FractionalIdealFunctions2
--R FRNAAF2  FramedNonAssociativeAlgebraFunctions2
--R FRUTIL   FactoredFunctionUtilities    FS2      FunctionSpaceFunctions2
--R FS2EXPXP FunctionSpaceToExponentialExpansion
--R FS2UPS   FunctionSpaceToUnivariatePowerSeries
--R FSAGG2   FiniteSetAggregateFunctions2
--R FSCINT   FunctionSpaceComplexIntegration
--R FSINT    FunctionSpaceIntegration     FSPECF   FunctionalSpecialFunction
--R FSPRMELT FunctionSpacePrimitiveElement
--R FSRED    FunctionSpaceReduce
--R FSUPFACT FunctionSpaceUnivariatePolynomialFactor
--R GALFACT  GaloisGroupFactorizer
--R GALFACTU GaloisGroupFactorizationUtilities
--R GALPOLYU GaloisGroupPolynomialUtilities
--R GALUTIL  GaloisGroupUtilities         GAUSSFAC GaussianFactorizationPackage
--R GB       GroebnerPackage
--R GBEUCLID EuclideanGroebnerBasisPackage
--R GBF      GroebnerFactorizationPackage GBINTERN GroebnerInternalPackage
--R GDRAW    GnuDraw                      GENEEZ   GenExEuclid
--R GENMFACT GeneralizedMultivariateFactorize
--R GENPGCD  GeneralPolynomialGcdPackage  GENUFACT GenUFactorize
--R GENUPS   GenerateUnivariatePowerSeries
--R GHENSEL  GeneralHenselPackage         GOSPER   GosperSummationMethod
--R GPAFF    GeneralPackageForAlgebraicFunctionField
--R GRAY     GrayCode                     GRDEF    GraphicsDefaults
--R GROEBSOL GroebnerSolve                GUESS    Guess
--R GUESSAN  GuessAlgebraicNumber         GUESSF   GuessFinite
--R GUESSF1  GuessFiniteFunctions         GUESSINT GuessInteger
--R GUESSP   GuessPolynomial              GUESSUP  GuessUnivariatePolynomial

--R HB       HallBasis                    HEUGCD   HeuGcd
--R IALGFACT InnerAlgFactor
--R IBACHIN  ChineseRemainderToolsForIntegralBases
--R IBATOOL  IntegralBasisTools           IBPTOOLS IntegralBasisPolynomialTools
--R ICDEN    InnerCommonDenominator       IDECOMP  IdealDecompositionPackage
--R IMATLIN  InnerMatrixLinearAlgebraFunctions
--R IMATQF   InnerMatrixQuotientFieldFunctions
--R INBFF    InnerNormalBasisFieldFunctions
--R INCRMAPS IncrementingMaps             INEP     InnerNumericEigenPackage
--R INFINITY Infinity                     INFORM1  InputFormFunctions1
--R INFPROD0 InfiniteProductCharacteristicZero
--R INFSP    InnerNumericFloatSolvePackage
--R INMODGCD InnerModularGcd              INNMFACT InnerMultFact
--R INPRODFF InfiniteProductFiniteField   INPRODPF InfiniteProductPrimeField
--R INPSIGN  InnerPolySign                INTAF    AlgebraicIntegration
--R INTALG   AlgebraicIntegrate           INTBIT   IntegerBits
--R INTDIVP  IntersectionDivisorPackage   INTEF    ElementaryIntegration
--R INTERGB  InterfaceGroebnerPackage     INTFACT  IntegerFactorizationPackage
--R INTFRSP  InterpolateFormsPackage      INTG0    GenusZeroIntegration
--R INTHEORY IntegerNumberTheoryFunctions INTHERAL AlgebraicHermiteIntegration
--R INTHERTR TranscendentalHermiteIntegration
--R INTPACK  AnnaNumericalIntegrationPackage
--R INTPAF   PureAlgebraicIntegration     INTPM    PatternMatchIntegration
--R INTRAT   RationalIntegration          INTRET   IntegerRetractions
--R INTRF    RationalFunctionIntegration
--R INTSLPE  IntegerSolveLinearPolynomialEquation
--R INTTOOLS IntegrationTools             INTTR    TranscendentalIntegration
--R INVLAPLA InverseLaplaceTransform      IPRNTPK  InternalPrintPackage
--R IR2      IntegrationResultFunctions2  IR2F     IntegrationResultToFunction
--R IROOT    IntegerRoots                 IRREDFFX IrredPolyOverFiniteField
--R IRRF2F   IntegrationResultRFToFunction
--R IRSN     IrrRepSymNatPackage
--R IRURPK   InternalRationalUnivariateRepresentationPackage
--R ISUMP    InnerPolySum                 ITFUN2   InfiniteTupleFunctions2
--R ITFUN3   InfiniteTupleFunctions3
--R ITRIGMNP InnerTrigonometricManipulations
--R KERNEL2  KernelFunctions2             KOVACIC  Kovacic
--R LAPLACE  LaplaceTransform             LAZM3PK  LazardSetSolvingPackage
--R LEADCDET LeadingCoefDetermination     LEXTRIPK LexTriangularPackage
--R LF       LiouvillianFunction          LGROBP   LinGroebnerPackage
--R LIMITPS  PowerSeriesLimitPackage      LIMITRF  RationalFunctionLimitPackage
--R LINDEP   LinearDependence             LIST2    ListFunctions2
--R LIST2MAP ListToMap                    LIST3    ListFunctions3
--R LISYSER  LinearSystemFromPowerSeriesPackage
--R LODEEF   ElementaryFunctionLODESolver
--R LODOF    LinearOrdinaryDifferentialOperatorFactorizer
--R LODOOPS  LinearOrdinaryDifferentialOperatorsOps
--R LOP      LinesOpPack
--R LPARSPT  LocalParametrizationOfSimplePointPackage
--R LPEFRAC  LinearPolynomialEquationByFractions

--R LSMP     LinearSystemMatrixPackage    LSMP1    LinearSystemMatrixPackage1
--R LSPP     LinearSystemPolynomialPackage
--R MAMA     MatrixManipulation           MAPHACK1 MappingPackageInternalHacks1
--R MAPHACK2 MappingPackageInternalHacks2 MAPHACK3 MappingPackageInternalHacks3
--R MAPPKG1  MappingPackage1              MAPPKG2  MappingPackage2
--R MAPPKG3  MappingPackage3              MAPPKG4  MappingPackage4
--R MATCAT2  MatrixCategoryFunctions2     MATLIN   MatrixLinearAlgebraFunctions
--R MATSTOR  StorageEfficientMatrixOperations
--R MCALCFN  MultiVariableCalculusFunctions
--R MCDEN    MatrixCommonDenominator
--R MDDFACT  ModularDistinctDegreeFactorizer
--R MESH     MeshCreationRoutinesForThreeDimensions
--R MFINFACT MultFiniteFactorize          MHROWRED ModularHermitianRowReduction
--R MKBCFUNC MakeBinaryCompiledFunction   MKFLCFN  MakeFloatCompiledFunction
--R MKFUNC   MakeFunction                 MKRECORD MakeRecord
--R MKUCFUNC MakeUnaryCompiledFunction    MLIFT    MultivariateLifting
--R MMAP     MultipleMap                  MONOTOOL MonomialExtensionTools
--R MPC2     MPolyCatFunctions2           MPC3     MPolyCatFunctions3
--R MPCPF    MPolyCatPolyFactorizer
--R MPRFF    MPolyCatRationalFunctionFactorizer
--R MRATFAC  MRationalFactorize           MRF2     MonoidRingFunctions2
--R MSYSCMD  MoreSystemCommands           MTHING   MergeThing
--R MULTFACT MultivariateFactorize        MULTSQFR MultivariateSquareFree
--R NAGC02   NagPolynomialRootsPackage    NAGC05   NagRootFindingPackage
--R NAGC06   NagSeriesSummationPackage    NAGD01   NagIntegrationPackage
--R NAGD02   NagOrdinaryDifferentialEquationsPackage
--R NAGD03   NagPartialDifferentialEquationsPackage
--R NAGE01   NagInterpolationPackage      NAGE02   NagFittingPackage
--R NAGE04   NagOptimisationPackage       NAGF01   NagMatrixOperationsPackage
--R NAGF02   NagEigenPackage
--R NAGF04   NagLinearEquationSolvingPackage
--R NAGF07   NagLapack                    NAGS     NagSpecialFunctionsPackage
--R NAGSP    NAGLinkSupportPackage        NCEP     NumericComplexEigenPackage
--R NCNTFRAC NumericContinuedFraction
--R NCODIV   NonCommutativeOperatorDivision
--R NEWTON   NewtonInterpolation          NFINTBAS NumberFieldIntegralBasis
--R NLINSOL  NonLinearSolvePackage        NODE1    NonLinearFirstOrderODESolver
--R NONE1    NoneFunctions1               NORMMA   NormInMonogenicAlgebra
--R NORMPK   NormalizationPackage         NORMRETR NormRetractPackage
--R NPCOEF   NPCoef                       NPOLYGON NewtonPolygon
--R NREP     NumericRealEigenPackage
--R NSUP2    NewSparseUnivariatePolynomialFunctions2
--R NTPOLFN  NumberTheoreticPolynomialFunctions
--R NUMERIC  Numeric                      NUMFMT   NumberFormats
--R NUMODE   NumericalOrdinaryDifferentialEquations
--R NUMQUAD  NumericalQuadrature          NUMTUBE  NumericTubePlot
--R OCTCT2   OctonionCategoryFunctions2   ODECONST ConstantLODE
--R ODEEF    ElementaryFunctionODESolver  ODEINT   ODEIntegration
--R ODEPACK  AnnaOrdinaryDifferentialEquationPackage
--R ODEPAL   PureAlgebraicLODE            ODEPRIM  PrimitiveRatDE

--R ODEPRRIC PrimitiveRatRicDE            ODERAT   RationalLODE
--R ODERED   ReduceLODE                   ODERTRIC RationalRicDE
--R ODESYS   SystemODESolver              ODETOOLS ODETools
--R OMEXPR   ExpressionToOpenMath         OMPKG    OpenMathPackage
--R OMSERVER OpenMathServerPackage        ONECOMP2 OnePointCompletionFunctions2
--R OPQUERY  OperationsQuery
--R OPTPACK  AnnaNumericalOptimizationPackage
--R ORDCOMP2 OrderedCompletionFunctions2  ORDFUNS  OrderingFunctions
--R OREPCTO  UnivariateSkewPolynomialCategoryOps
--R ORTHPOL  OrthogonalPolynomialFunctions
--R OUT      OutputPackage                PADE     PadeApproximants
--R PADEPAC  PadeApproximantPackage
--R PAFF     PackageForAlgebraicFunctionField
--R PAFFFF   PackageForAlgebraicFunctionFieldOverFiniteField
--R PAN2EXPR PolynomialAN2Expression      PARAMP   ParametrizationPackage
--R PARPC2   ParametricPlaneCurveFunctions2
--R PARSC2   ParametricSpaceCurveFunctions2
--R PARSU2   ParametricSurfaceFunctions2  PARTPERM PartitionsAndPermutations
--R PATMATCH PatternMatch                 PATRES2  PatternMatchResultFunctions2
--R PATTERN1 PatternFunctions1            PATTERN2 PatternFunctions2
--R PCOMP    PolynomialComposition        PDECOMP  PolynomialDecomposition
--R PDEPACK  AnnaPartialDifferentialEquationPackage
--R PERMAN   Permanent
--R PFBR     PolynomialFactorizationByRecursion
--R PFBRU    PolynomialFactorizationByRecursionUnivariate
--R PFO      PointsOfFiniteOrder          PFOQ     PointsOfFiniteOrderRational
--R PFORP    PackageForPoly               PFOTOOLS PointsOfFiniteOrderTools
--R PFRPAC   PartialFractionPackage       PGCD     PolynomialGcdPackage
--R PGE      PermutationGroupExamples     PGROEB   PolyGroebner
--R PICOERCE PiCoercions                  PINTERP  PolynomialInterpolation
--R PINTERPA PolynomialInterpolationAlgorithms
--R PLEQN    ParametricLinearEquations    PLOT1    PlotFunctions1
--R PLOTTOOL PlotTools                    PLPKCRV  PolynomialPackageForCurve
--R PMASS    PatternMatchAssertions       PMASSFS  FunctionSpaceAssertions
--R PMDOWN   PatternMatchPushDown         PMFS     PatternMatchFunctionSpace
--R PMINS    PatternMatchIntegerNumberSystem
--R PMKERNEL PatternMatchKernel           PMLSAGG  PatternMatchListAggregate
--R PMPLCAT  PatternMatchPolynomialCategory
--R PMPRED   AttachPredicates
--R PMPREDFS FunctionSpaceAttachPredicates
--R PMQFCAT  PatternMatchQuotientFieldCategory
--R PMSYM    PatternMatchSymbol           PMTOOLS  PatternMatchTools
--R PNTHEORY PolynomialNumberTheoryFunctions
--R POLTOPOL PolToPol
--R POLUTIL  RealPolynomialUtilitiesPackage
--R POLY2    PolynomialFunctions2
--R POLY2UP  PolynomialToUnivariatePolynomial
--R POLYCATQ PolynomialCategoryQuotientFunctions
--R POLYLIFT PolynomialCategoryLifting    POLYROOT PolynomialRoots
--R POLYVEC  U32VectorPolynomialOperations

--R PREASSOC PrecomputedAssociatedEquations
--R PRIMARR2 PrimitiveArrayFunctions2     PRIMELT  PrimitiveElement
--R PRIMES   IntegerPrimesPackage         PRINT    PrintPackage
--R PRJALGPK ProjectiveAlgebraicSetPackage
--R PRS      PseudoRemainderSequence
--R PSETPK   PolynomialSetUtilitiesPackage
--R PSEUDLIN PseudoLinearNormalForm       PSQFR    PolynomialSquareFree
--R PTFUNC2  PointFunctions2              PTPACK   PointPackage
--R PUSHVAR  PushVariables
--R PWFFINTB PAdicWildFunctionFieldIntegralBasis
--R QALGSET2 QuasiAlgebraicSet2           QCMPACK  QuasiComponentPackage
--R QFCAT2   QuotientFieldCategoryFunctions2
--R QUATCT2  QuaternionCategoryFunctions2 RADUTIL  RadixUtilities
--R RANDSRC  RandomNumberSource           RATFACT  RationalFactorize
--R RATRET   RationalRetractions          RDEEF    ElementaryRischDE
--R RDEEFS   ElementaryRischDESystem      RDETR    TranscendentalRischDE
--R RDETRS   TranscendentalRischDESystem  RDIST    RandomDistributions
--R RDIV     ReducedDivisor               REAL0    RealZeroPackage
--R REAL0Q   RealZeroPackageQ             REALSOLV RealSolvePackage
--R RECOP    RecurrenceOperator           REDORDER ReductionOfOrder
--R REP      RadicalEigenPackage          REP1     RepresentationPackage1
--R REP2     RepresentationPackage2       REPDB    RepeatedDoubling
--R REPSQ    RepeatedSquaring             RESLATC  ResolveLatticeCompletion
--R RETSOL   RetractSolvePackage          RF       RationalFunction
--R RFDIST   RandomFloatDistributions     RFFACT   RationalFunctionFactor
--R RFFACTOR RationalFunctionFactorizer   RFP      RootsFindingPackage
--R RIDIST   RandomIntegerDistributions   RINTERP  RationalInterpolation
--R RMCAT2   RectangularMatrixCategoryFunctions2
--R RSDCMPK  RegularSetDecompositionPackage
--R RSETGCD  RegularTriangularSetGcdPackage
--R RURPK    RationalUnivariateRepresentationPackage
--R SAEFACT  SimpleAlgebraicExtensionAlgFactor
--R SAERFFC  SAERationalFunctionAlgFactor SCACHE   SortedCache
--R SCPKG    StructuralConstantsPackage   SEG2     SegmentFunctions2
--R SEGBIND2 SegmentBindingFunctions2
--R SFQCMPK  SquareFreeQuasiComponentPackage
--R SFRGCD   SquareFreeRegularTriangularSetGcdPackage
--R SGCF     SymmetricGroupCombinatoricFunctions
--R SHP      SturmHabichtPackage          SIGNEF   ElementaryFunctionSign
--R SIGNRF   RationalFunctionSign
--R SIMPAN   SimplifyAlgebraicNumberConvertPackage
--R SMITH    SmithNormalForm              SOLVEFOR PolynomialSolveByFormulas
--R SOLVERAD RadicalSolvePackage          SOLVESER TransSolvePackageService
--R SOLVETRA TransSolvePackage            SORTPAK  SortPackage
--R SPECOUT  SpecialOutputPackage
--R SRDCMPK  SquareFreeRegularSetDecompositionPackage
--R STINPROD StreamInfiniteProduct        STNSR    StreamTensor
--R STREAM1  StreamFunctions1             STREAM2  StreamFunctions2
--R STREAM3  StreamFunctions3             STTAYLOR StreamTaylorSeriesOperations
--R STTF     StreamTranscendentalFunctions

--R STTFNC   StreamTranscendentalFunctionsNonCommutative
--R SUBRESP  SubResultantPackage          SUMFS    FunctionSpaceSum
--R SUMRF    RationalFunctionSum
--R SUP2     SparseUnivariatePolynomialFunctions2
--R SUPFRACF SupFractionFactorizer        SYMFUNC  SymmetricFunctions
--R SYSSOLP  SystemSolvePackage           TABLBUMP TableauxBumpers
--R TANEXP   TangentExpansions            TBCMPPK  TabulatedComputationPackage
--R TEMUTL   TemplateUtilities            TEX1     TexFormat1
--R TOOLSIGN ToolsForSign                 TOPSP    TopLevelThreeSpace
--R TRIGMNIP TrigonometricManipulations   TRIMAT   TriangularMatrixOperations
--R TRMANIP  TranscendentalManipulations  TUBETOOL TubePlotTools
--R TWOFACT  TwoFactorize                 UDPO     UserDefinedPartialOrdering
--R UDVO     UserDefinedVariableOrdering
--R UFPS1    UnivariateFormalPowerSeriesFunctions
--R ULS2     UnivariateLaurentSeriesFunctions2
--R UNIFACT  UnivariateFactorize          UNISEG2  UniversalSegmentFunctions2
--R UP2      UnivariatePolynomialFunctions2
--R UPCDEN   UnivariatePolynomialCommonDenominator
--R UPDECOMP UnivariatePolynomialDecompositionPackage
--R UPDIVP   UnivariatePolynomialDivisionPackage
--R UPMP     UnivariatePolynomialMultiplicationPackage
--R UPOLYC2  UnivariatePolynomialCategoryFunctions2
--R UPSQFREE UnivariatePolynomialSquareFree
--R UPXS2    UnivariatePuiseuxSeriesFunctions2
--R UTS2     UnivariateTaylorSeriesFunctions2
--R UTSODE   UnivariateTaylorSeriesODESolver
--R UTSODETL UTSodetools                  UTSSOL   TaylorSolve
--R VECTOR2  VectorFunctions2             VIEW     ViewportPackage
--R VIEWDEF  ViewDefaultsPackage          WEIER    WeierstrassPreparation
--R WFFINTBS WildFunctionFieldIntegralBasis
--R XEXPPKG  XExponentialPackage
--R YSTREAM  ParadoxicalCombinatorsForStreams
--R ZDSOLVE  ZeroDimensionalSolvePackage  ZLINDEP  IntegerLinearDependence
--E 15

———-

This exercises setOutputLibrary, setInputLibrary, and setAsharpArgs — * —

--S 16 of 20
)set compiler
--R
--R                  Current Values of  compiler  Variables
--R
--RVariable     Description                                Current Value
--R-----------------------------------------------------------------------------
--Routput       library in which to place compiled code    user.lib
--Rinput        controls libraries from which to load compiled code
--R

--E 16

———-

This exercises setInputLibrary, describeInputLibraryArgs — * —

--S 17 of 20
)set compiler input
--R
--R---------------------------- The input Option -----------------------------
--R
--R Description: controls libraries from which to load compiled code
--R
--R )set compile input add library is used to tell AXIOM to add library to
--Rthe front of the path used to find compile code.
--R )set compile input drop library is used to tell AXIOM to remove library
--Rfrom this path.
--E 17
--S 18 of 20
)set compiler input add
--R
--R )set compile input add library is used to tell AXIOM to add library to
--Rthe front of the path used to find compile code.
--R )set compile input drop library is used to tell AXIOM to remove library
--Rfrom this path.
--E 18
--S 19 of 20
)set compiler input add foo
--E 19

———-

This exercises setOutputLibrary — * —

--S 20 of 20
)set compiler output
--R
--R---------------------------- The output Option ----------------------------
--R
--R Description: library in which to place compiled code
--R
--R )set compile output library is used to tell the compiler where to place
--Rcompiled code generated by the library compiler.  By default it goes
--Rin a file called user.lib in the current directory.
--E 20
)spool

)lisp (bye) ———-

22

References^{[1] nothing}

23